A coin is tossed 225 times. Estimate the chance of getting 115 heads and 110 tails.
Answer by stanbon(75887)

stanford university accepted 2300 of 32000 freshmen applicants for fall of 2010. Find the probability that none of 20 randomly selected applicants were accepted.
------------
P(accepted) = 2300/32000
----
P(20 selected) = (2300/32000)^20 = 1.35x10^-23
===============
Cheers,
Stan H.